OpenStudy (anonymous):

Anyone will fane and medal Solve the following system: y = x + 3 3x + y = 19 (7, 4) (−4, 7) (4, 7) (4, −7)

OpenStudy (abb0t):

What's "fane" ?

OpenStudy (anonymous):

@jjamz87 ? have a clue

OpenStudy (anonymous):

its a typo **** i meant fan

OpenStudy (anonymous):

@abb0t

OpenStudy (jjamz87):

Ok. I'll explain it as well.

OpenStudy (anonymous):

please i really wanna understand , not just get the answer ^^ (:

pooja195 (pooja195):

Plug in equation 1 into equation 2 3x + (x+3) = 19 now solve for x

pooja195 (pooja195):

Combine like terms :)

OpenStudy (anonymous):

oh okay so it would be 4x+3=19

OpenStudy (abb0t):

You can rearrange, using algebra tricks, to manipulate an substitute. That means you can rearrange equation 1: \(\sf \color{red}{y-3 = x}\) notice it says x equals y - 3? well, yeah, thats what it equals to, so substitute THAT thing that x equals to into equation 2: \(\sf \color{blue}{3}(\color{red}{y-3})\color{blue}{+y=19}\)

pooja195 (pooja195):

right subract 3 from both sides 4x=19-3

OpenStudy (jjamz87):

y = x + 3 3x + y =19 plug in the first equation into y 3x + (x + 3) = 19 simplify 3x + x + 3 = 19 4x = 16 divide 4 from both sides. x = 4

OpenStudy (anonymous):

4x=16

pooja195 (pooja195):

whats 16/4 ?

OpenStudy (abb0t):

4

OpenStudy (anonymous):

4

pooja195 (pooja195):

good therfore x=4 now plug that into the 1st equation y=4+3 whats 4+3?

OpenStudy (anonymous):

okay so i just have to rearrange @abb0t for these questions and plug it in to the other equation

OpenStudy (anonymous):

thank you @jjamz87

pooja195 (pooja195):

@awlas15 rearraging makes it more complicated you can leave it as is...

pooja195 (pooja195):

*rearranging

OpenStudy (anonymous):

ohh okay so then the answer 4,7

pooja195 (pooja195):

yep

OpenStudy (abb0t):

No. Do not listen to her, rearranging is much easier.

OpenStudy (abb0t):

It is simply plug and chug and combining like terms.

OpenStudy (jjamz87):

@awlas15 yes you are correct

OpenStudy (abb0t):

Nice one @jjamz87

pooja195 (pooja195):

@abb0t quit trying to make things more harder than they need to be.

OpenStudy (abb0t):

Good job, @awlas15

OpenStudy (abb0t):

If you need any more answers, tag me :)

OpenStudy (jjamz87):

Y'all stop fighting I have already answered the question.

OpenStudy (anonymous):

thanks everyone !!! how can i give you guys all medals ): and i guess harder is more better to @abb0t perspective

OpenStudy (abb0t):

You did a fabulous job, @awlas15

pooja195 (pooja195):

Giving the answer is not necessarily helping. But ok. Whatever.

OpenStudy (jjamz87):

Excuse me but I explained it.

OpenStudy (abb0t):

Sounds like someone is mad that they earned the medal and they didn't. Ouchies! Some people are better at explaining than others :)

pooja195 (pooja195):

LOL i dont care about medals

pooja195 (pooja195):

Anyways if a mod does nothing why should i? Aight have a nice day.

OpenStudy (anonymous):

seriously guys thanks i mean its cause i asked him first but if anything ill give you a medal on any other question I'm sorry @pooja195 but seriously thank you guys and quite fighting i understood it thats the point of open study @pooja195 @jjamz87 @abb0t but AGAIN I THANK YOU I REALLY DO APPRECIATE IT LIKE SERIOUSLY THANKS

pooja195 (pooja195):

@awlas15 i dont care about medals

OpenStudy (anonymous):

I know but i feel bad i wanna give you a medal as well i hate how i can not but really thank you seriously mean a lot @pooja195